OpenStudy (snitchseeker1496):

the two legs of a triangle are 300 and 150 m each, respectively. The angle opposite the 150 m side is 26 degree. What is the 3rd side?

OpenStudy (wolf1728):

Here's a graphic:

OpenStudy (wolf1728):

And here's the Law of Sines:

OpenStudy (snitchseeker1496):

I got 341.78 m, am I right?

OpenStudy (wolf1728):

150/sin(26) = 300/ angle Yes, this will just give you the second angle

OpenStudy (snitchseeker1496):



OpenStudy (wolf1728):

snitch - I get 341.78 too

OpenStudy (snitchseeker1496):



OpenStudy (snitchseeker1496):

Thank you guys :D

OpenStudy (wolf1728):

Since this problem gives you side side angle, there can be another solution 197.49

OpenStudy (snitchseeker1496):

can you teach me for the other alternative solution? :)

jimthompson5910 (jim_thompson5910):

This is one of the cases of SSA where you'll get 2 possible triangles See attached Triangle ABC is one possible triangle. Triangle ABD is another possible triangle.

OpenStudy (snitchseeker1496):

Thank you @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

forgot to label BC. I used geogebra to make the triangles

OpenStudy (wolf1728):

Wow - I just made a graphic, but Jim's looks really good.

jimthompson5910 (jim_thompson5910):

I had geogebra do all of the work @wolf1728 https://www.geogebra.org/

OpenStudy (snitchseeker1496):

Can I download this for free?

jimthompson5910 (jim_thompson5910):

yes you can also use it as a web app too, which means it runs without you having to download anything. I prefer the desktop version though

OpenStudy (snitchseeker1496):

Thanks again guys :)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (wolf1728):

150/sin(26 degrees) = 300/ sin(Angle C) sin(Angle C) = sin(26) * 2 sin(Angle C) = 0.43837 * 2 sin(Angle C) = 0.87674 Angle C = 118.748

OpenStudy (wolf1728):

Angle B = 35.252

OpenStudy (wolf1728):

Law of Cosines b^2 = a^2+c^2 -2*a*c * cos(B) b^2 = 150^2 + 300^2 -2*150*300 * cos(35.252) b^2 = 112,500 -90,000 * 0.81662 b^2 = 112,500 -73,496 b^2 = 39004 b = 197.49

OpenStudy (snitchseeker1496):

the answer should be 197.49?

OpenStudy (snitchseeker1496):

thank you :D